### v2

@@ 4.2 Model Validation
One key challenge in these environments is that the consumption distribution follows a heavy-tail distribution, with some goods having an outsized share of consumption (McKenzie, 2023), and the challenge with traditional characteristics space models is that the good characteristics such as cast and genre cannot rationalize these patterns. Thus, our first simple validation exercise, presented in Figure 5, shows that our model (in-sample) recovers good-level shares reasonably well.
@@ 4.2.2 Comparison to Experimental Estimates
We denote each of the treatment arms as $g\in\mathcal{G}$ and denote $P_{k}(g)$ as the probability that good $k$ gets chosen under the recommendation restriction $g$. Therefore, following Conlon and Mortimer (2021) we can use the Wald Estimator to obtain the main diversion measure of interest:
@@ 5.1.1 Value of the Current RecSys
Since we have separately isolated the effect of recommendations from the baseline preferences of users, we can credibly simulate the impact of different recommender systems on choices. We consider the impact of these different recommendation policies on two key variables: (i) overall engagement – how much incremental consumption is generated – and (ii) the diversity of goods that get consumed. Although overall engagement is the most relevant measure for the platform as it provides a measure for how much the recommender system is helping users find relevant goods, diversity is also important. This is because consumption diversity is beneficial in the long-run: it has been shown to correlate with higher long-run satisfaction for users (Anderson et al., 2020; Chen et al., 2024) and it helps to sustain a healthier two-sided market by distributing viewership more evenly across good producers.
@@
    Matrix Factorization Recommendations - $\mathcal{C}_{it}^{\text{FACTOR}}$: The recommendations are computed using a matrix factorization algorithm (Koren et al., 2009). This captures the “state of the art” approach to computing recommendations from 2015 as discussed in Gomez-Uribe and Hunt (2015). [^5] We employ a vanilla SVD matrix factorization approach, without tuning or hyperparameter optimization. This is intended to provide a reasonable baseline for personalized recommendations, but does not represent a specific recommendation system previously employed by Netflix.
To perform these counterfactual estimates, we simulate each alternative recommendation system by replacing the set of recommended goods according to the relevant algorithm, keeping the number of recommendations each user receives constant. Our implementation takes into account several complexities of the actual good interface and business logic. Specifically, we model the effects of recommendations via a separate parameter for each of four different locations on the page the recommendations are present. This accounts for the fact that recommendations in certain locations (e.g., as described in Section 2) are more salient and influential. When constructing counterfactual recommendations, we also disallow recommending goods that a user has already watched, in line with the platform’s discovery-oriented business rules. Changes to these modeling choices can significantly affect the resulting engagement and diversity metrics for a counterfactual recommendation algorithm, but not the relative ordering.
@@
First, these results imply that the current algorithm leads to large improvements in engagement relative to each of the benchmark algorithms. Reverting to random, popularity, or matrix factorization algorithms would reduce engagement by 16%, 12%, and 4%, respectively. These gains are quantitatively large and economically meaningful—e.g., internal estimates would imply that reverting back to matrix factorization recommendations would lead to similar monetary losses as the estimates from Gomez-Uribe and Hunt (2015) indicate reverting back to popularity from matrix factorization would induce in 2015.666These estimates are likely lower-bounds, given that our implementation of the recommendation bonus implicitly incorporates the trust that the users have with the RecSys, which would likely be lower under these counterfactual policies. This improvement highlights that, despite recent concerns in the RecSys community that modern algorithmic improvements are providing little gain relative to traditional approaches such as matrix factorization (Ferrari Dacrema et al., 2019), the latest algorithmic improvements applied at one of the most prominent deployments of RecSys are leading to meaningful gains.
@@
Overall, these results show that the current RecSys meaningfully improves engagement while maintaining relatively high consumption diversity. In contrast, traditional matrix-factorization and popularity-based algorithms would lead to large reductions in engagement and concentrate viewership on a much narrower set of goods. This underscores that, despite the concerns in Ferrari Dacrema et al. (2019), modern improvements to recommendation algorithms are both helping match users with their preferred goods and resulting in improved consumption diversity which is important for sustaining a broader good ecosystem.
@@ 5.1.2 Good-Level Value of Recommendation
Thus, using our model we can simulate both potential outcomes. To estimate $Y_{ij}(1)$ for each user $i$, we force good $j$ to be in each user’s recommended set [^7] If $j\in\mathcal{C}_{it}$ originally, then we leave the rest of $\mathcal{C}_{it}$ same. If $j\notin\mathcal{C}_{it}$, then we take the average consumption probability of NN random goods to be replaced with good $j$ in $\mathcal{C}_{it}$. and to estimate $Y_{ij}(0)$ we remove good $j$ from $\mathcal{C}_{it}$. Thus, given this the causal effect of recommending good $j$ to user $i$ is $\tau_{ij}\equiv Y_{ij}(1)-Y_{ij}(0)$.
@@ 5.2 The Incremental Demand for Goods
For the first challenge, we can rely on the utility-based approach for imputing counterfactual recommendations – described in Appendix A – that uses draws from the predicted user-specific utilities. [^9] We rely on the utility-based estimate procedure since changing the catalog would mean that the recommendation model is based on recommendations under a different catalog which would likely produce incorrect recommendations, especially for new goods. For the second challenge, we provide an extension of the model that relies on an “exogenous” characteristic space that allows us to simulate counterfactual engagement for new goods. [^10] This has historically been one of the most prominent advantages of characteristic-space approaches to demand (Petrin, 2002).
Model Extension: Exogenous Good Embeddings. We therefore consider an extension of our model that replaces the endogenously learned good-level embeddings with an exogenous set of good characteristics. This allows us to credibly estimate counterfactual engagement for new goods. The empirical challenge is in constructing these good characteristics as this historically has prevented reasonable models of movie demand. We rely on good-level embeddings developed internally at Netflix that provide a high-dimensional representation of a good and capture many hard to directly measure characteristics. [^11] Recent examples in economics that rely on embeddings in lieu of traditional characteristics spaces in demand models include Quan and Williams (2019); Toubia et al. (2019); Chen et al. (2020); Bach et al. (2024); Magnolfi et al. (2025); Compiani et al. (2025). Importantly, these embeddings are trained based on pre-consumption characteristics and are primarily driven by human tags associated with each of the goods that are combined with observable characteristics, such as cast and genre, to produce a high-dimensional representation of both existing goods and new goods that are not available on the platform.
@@ 6 Conclusion
Our paper highlights the central role of recommendation in driving user choices in these markets and the large value of the current RecSys on Netflix in several ways. First, exploration experiments in the RecSys can be a powerful lever to learn substitution patterns across goods. Second, we show that compared to both non-personalized and previous state-of-the-art methods for recommendations the current RecSys drives a meaningfully large amount of additional engagement. Third, we decompose the effects of recommendation on individual goods into selection into who gets targeted, the exposure effect of a recommendation, and differential targeting of the recommendations. We show that while exposure meaningfully increases consumption, effective targeting plays a quantitatively larger role and that this effect is most pronounced for mid-sized goods, as opposed to very niche or broadly appealing goods. Overall, our paper provides rich insights into how to measure the value of recommendations for the platform and their impact on consumer behavior.
@@ Appendix A Imputing Counterfactual Recommendations
The utility-based procedure relies on the intuition that goods that appear in recommendations are those that have high predicted utility for user ii. Thus, if a good is not allowed to appear in the recommendations, the next most substitutable good will also be one that has high predicted utility. While, of course, there is much more nuance in practice to the recommendation system, our procedure follows this intuition by sampling goods using a softmax rule based on their predicted utility. Specifically, we consider the following replacement rule for $j\notin\mathcal{C}_{it}$:

### v3

@@ 4.2 Model Validation
One key challenge in these environments is that the consumption distribution follows a heavy-tail distribution, with some goods having an outsized share of consumption (mckenzie2023economics), and the challenge with traditional characteristics space models is that the good characteristics such as cast and genre cannot rationalize these patterns. Thus, our first simple validation exercise, presented in Figure 5, shows that our model (in-sample) recovers good-level shares reasonably well.
@@ 4.2.2 Comparison to Experimental Estimates
We denote each of the treatment arms as $g\in\mathcal{G}$ and denote $P_{k}(g)$ as the probability that good $k$ gets chosen under the recommendation restriction $g$. Therefore, following conlon2021empirical we can use the Wald Estimator to obtain the main diversion measure of interest:
@@ 5.1.1 Value of the Current RecSys
Since we have separately isolated the effect of recommendations from the baseline preferences of users, we can credibly simulate the impact of different recommender systems on choices. We consider the impact of these different recommendation policies on two key variables: (i) overall engagement – how much incremental consumption is generated – and (ii) the diversity of goods that get consumed. Although overall engagement is the most relevant measure for the platform as it provides a measure for how much the recommender system is helping users find relevant goods, diversity is also important. This is because consumption diversity is beneficial in the long-run: it enables users to benefit from a broader set of goods and has been shown to correlate with higher long-run satisfaction (anderson2020algorithmic; chen2024impact).
@@
    Matrix Factorization Recommendations - $\mathcal{C}_{it}^{\text{FACTOR}}$: The recommendations are computed using a matrix factorization algorithm (koren2009matrix). This captures the “state of the art” approach to computing recommendations from 2015 as discussed in gomez2015netflix. [^5] We employ a vanilla SVD matrix factorization approach, without tuning or hyperparameter optimization. This is intended to provide a reasonable baseline for personalized recommendations, but does not represent a specific recommendation system previously employed by Netflix.
To perform these counterfactual estimates, we simulate each alternative recommendation system by replacing the set of recommended goods according to the relevant algorithm, keeping the number of recommendations each user receives constant. Our implementation takes into account several complexities of the actual good interface and business logic. Specifically, we model the effects of recommendations via a separate parameter for each of four different locations on the page the recommendations are present. This accounts for the fact that recommendations in certain locations (e.g., as described in Section 2) are more salient and influential. When constructing counterfactual recommendations, we also disallow recommending goods that a user has already watched, as an approximation of the platform’s discovery-oriented recommendations. Changes to these modeling choices can significantly affect the resulting engagement and diversity metrics for a counterfactual recommendation algorithm, but not the relative ordering.
@@
First, these results imply that the current algorithm leads to large improvements in engagement relative to each of the benchmark algorithms. Reverting to random, popularity, or matrix factorization algorithms would reduce engagement by 16%, 12%, and 4%, respectively. These gains are quantitatively large and economically meaningful—e.g., internal estimates would imply that reverting back to matrix factorization recommendations would lead to similar monetary losses as the estimates from gomez2015netflix indicate reverting back to popularity from matrix factorization would induce in 2015.666These estimates are likely lower-bounds, given that our implementation of the recommendation bonus implicitly incorporates the trust that the users have with the RecSys, which would likely be lower under these counterfactual policies. This improvement highlights that, despite recent concerns in the RecSys community that modern algorithmic improvements are providing little gain relative to traditional approaches such as matrix factorization (ferrari2019we), the latest algorithmic improvements applied at one of the most prominent deployments of RecSys are leading to meaningful gains.
@@
Overall, these results show that the current RecSys meaningfully improves engagement while maintaining relatively high consumption diversity. In contrast, traditional matrix-factorization and popularity-based algorithms would lead to large reductions in engagement and concentrate viewership on a much narrower set of goods. This underscores that, despite the concerns in ferrari2019we, modern improvements to recommendation algorithms are both helping match users with their preferred goods and resulting in improved consumption diversity which is important for sustaining a broader good ecosystem.
@@ 5.1.2 Good-Level Value of Recommendation
Thus, using our model we can simulate both potential outcomes. To estimate $Y_{ij}(1)$ for each user $i$, we force good $j$ to be in each user’s recommended set [^7] If $j\in\mathcal{C}_{it}$ originally, then we leave the rest of $\mathcal{C}_{it}$ same. If $j\notin\mathcal{C}_{it}$, then we take the average consumption probability of $N$ random goods to be replaced with good $j$ in $\mathcal{C}_{it}$. and to estimate $Y_{ij}(0)$ we remove good $j$ from $\mathcal{C}_{it}$. Thus, given this the causal effect of recommending good $j$ to user $i$ is $\tau_{ij}\equiv Y_{ij}(1)-Y_{ij}(0)$.
@@ 5.2 The Incremental Demand for Goods
For the first challenge, we can rely on the utility-based approach for imputing counterfactual recommendations – described in Appendix A – that uses draws from the predicted user-specific utilities. [^9] We rely on the utility-based estimate procedure since changing the catalog would mean that the recommendation model is based on recommendations under a different catalog which would likely produce incorrect recommendations, especially for new goods. For the second challenge, we provide an extension of the model that relies on an “exogenous” characteristic space that allows us to simulate counterfactual engagement for new goods. [^10] This has historically been one of the most prominent advantages of characteristic-space approaches to demand (petrin2002quantifying).
Model Extension: Exogenous Good Embeddings. We therefore consider an extension of our model that replaces the endogenously learned good-level embeddings with an exogenous set of good characteristics. This allows us to credibly estimate counterfactual engagement for new goods. The empirical challenge is in constructing these good characteristics as this historically has prevented reasonable models of movie demand. We rely on good-level embeddings developed internally at Netflix that provide a high-dimensional representation of a good and capture many hard to directly measure characteristics. [^11] Recent examples in economics that rely on embeddings in lieu of traditional characteristics spaces in demand models include quan2019extracting; toubia2019extracting; chen2020studying; bach2024adventures; magnolfi2025triplet; compiani2025demand. Importantly, these embeddings are trained based on pre-consumption characteristics and are primarily driven by human tags associated with each of the goods that are combined with observable characteristics, such as cast and genre, to produce a high-dimensional representation of both existing goods and new goods that are not available on the platform.
@@ 6 Conclusion
Our paper highlights the central role of recommendation in driving user choices in these markets and the large value of the current RecSys on Netflix in several ways. First, exploration experiments in the RecSys can be a powerful lever to learn substitution patterns across goods. Second, we show that compared to both non-personalized and previous state-of-the-art methods for recommendations the current RecSys drives a meaningfully large amount of additional engagement. Third, we decompose the effects of recommendation on individual goods into selection into who gets targeted, the exposure effect of a recommendation, and differential responsiveness of targeted users from the recommendations. We show that while exposure meaningfully increases consumption, effective targeting plays a quantitatively larger role and that this effect is most pronounced for mid-sized goods, as opposed to very niche or broadly appealing goods. Overall, our paper provides rich insights into how to measure the value of recommendations for platforms and their impact on consumer behavior.
@@ Appendix A Imputing Counterfactual Recommendations
The utility-based procedure relies on the intuition that goods that appear in recommendations are those that have high predicted utility for user $i$. Thus, if a good is not allowed to appear in the recommendations, the next most substitutable good will also be one that has high predicted utility. While, of course, there is much more nuance in practice to the recommendation system, our procedure follows this intuition by sampling goods using a softmax rule based on their predicted utility. Specifically, we consider the following replacement rule for $j\notin\mathcal{C}_{it}$: